OpenStudy (megannicole51):

does the series converge or diverge? n!/(n+2)!

OpenStudy (megannicole51):

An+1=(n+1)!/(n+3)! An=n!/(n+2)!

OpenStudy (megannicole51):

when you cross multiply you get... (n+1)!/(n+3)!/(n+2)!/n!

OpenStudy (anonymous):

(n+1)! = (n+1)(n!) (n+3)! = (n+3)(n+2)!

OpenStudy (megannicole51):

then to get them to cancel out you have to use n(n-1)! on (n+3)!

hartnn (hartnn):

same thing we used in last question!

OpenStudy (megannicole51):

so then you end up getting (n+1)/(n+3)

OpenStudy (megannicole51):

which diverges

hartnn (hartnn):

i don't think ratio test will work here, you'll get limit = 1

OpenStudy (megannicole51):

how did u get 1?

hartnn (hartnn):

lim n-> infinity (n+1)/(n+3) = ...

OpenStudy (anonymous):

Long run behavior shows that it will be lim=1

OpenStudy (megannicole51):

oh crap...so then what test should i use?

OpenStudy (anonymous):

n!/(n+2)! = 1/[(n+2)(n+1)]

hartnn (hartnn):

so we are not using ratio test, right ?

OpenStudy (anonymous):

Correct, no ratio test needed.

OpenStudy (anonymous):

1/[(n+2)(n+1)] tends to 0, as n----> infinity.

OpenStudy (anonymous):

as the series is basically 1/n^2

OpenStudy (megannicole51):

how do u know its 0?

OpenStudy (anonymous):

1/n^2 is basically 1/very large number...as the denominator gets very large, 1/that very large number tends to 0.

OpenStudy (megannicole51):

i see it now! thank you!

OpenStudy (anonymous):

welcome.